OpenStudy (anonymous):

What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0? slope = -2, y-intercept = slope = 2, y-intercept = slope = -2, y-intercept = slope = 2, y-intercept = slope = -2, y-intercept =

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

ok.....to get rid of the parenthesis, you multiply the number outside of the parenthesis by everything in the parenthesis. For example : 2(x + 2) is the same as 2 * x + 2 * 2 = 2x + 4 do you understand that ? That is what's called " distributing through the parenthesis. are you with me so far ?

OpenStudy (anonymous):

A little bit

OpenStudy (anonymous):

part of your problem is 3(y - 2)... that is the same as : 3 * y + 3 * -2 = 3y + (-6) or just 3y - 6 understanding so far ?

OpenStudy (anonymous):

Little bit lol but go ahead

OpenStudy (anonymous):

so far, we have : 3y - 6 + 6(x + 1) - 2 = 0 now we will distribute the 6 through the other parenthesis 3y - 6 + 6x + 6 - 2 = 0 --- combine like terms 3y + 6x - 2 = 0 --- notice that the 6's cancel each other out 3y + 6x = 2 put in y = mx + b form 3y = -6x + 2 y = -2x + 2/3 now in y = mx + b form, m is your slope....so your slope is -2 and in y = mx + b form, the y int is in the b position.....so your y int is 2/3 any questions ?

OpenStudy (anonymous):

Nope! I understand now. Thank you!

OpenStudy (anonymous):

glad to help :)